# Concatenation of Kerr solitary waves in Ceramic YAG: application to coherent Raman imaging

### v4

@@ Abstract
A coherent concatenation of multiple Townes solitons may lead to a stable infrared and visible broadband filament in ceramic YAG polycrystal. This self-trapped soliton train is leveraged to implement self-referenced multiplex coherent anti-Stokes Raman scattering imaging. Simulations and experiments illustrating the filamentation process and the concatenation of focusing-defocusing cycles in ceramic and crystal YAG are presented. In addition, our simulations and experiments further examine the dependence of the filamentation onset location and supercontinuum generation upon peak input power. Understanding this dependence is key for implementation of viable CARS imaging techniques, due to the comparatively exceptional ability of YAG to generate supercontinuum which can enable higher-sensitivity imaging without delay lines.
Introduction Spatial solitons are formed by the deformation-free propagation of a wave packet, confined by the nonlinear effect of its own intensity [1]. Nonlinear saturation allows for wave stabilization over several Fresnel lengths, leading to a modulation-free, axially-symmetric beam. These so-called Townes solitons can balance diffraction via nonlinear effects at a given peak power. In two or more spatial dimensions, additional defocusing effects, such as plasma defocusing and nonlinear absorption, are known to arrest “catastrophic collapse” [1, 2, 3]. This process gives, at best, quasi-stability and can otherwise lead to cycles of focusing-defocusing events [3, 4, 5, 6, 7, 8]. It is also known that the extreme confinement of the field produces conical waves, causing some of the beam energy to propagate in a cone shape towards the center and feed the core filament due to a coherent interference effect [1, 5, 8].
In this work, we show experimentally and numerically that the self-focusing process obtained in a Nd:YAG ceramic rod can give rise to coherent, concatenated Townes solitons, producing a stable filament in both visible and infrared spectra. This stabilization is obtained thanks to the particular dielectric properties of ceramics, which facilitate emission of conical waves, plasma defocusing, and nonlinear/plasma absorption [5, 9]. Such broadband supercontinuum (SC) is surprisingly stable, and can be efficiently leveraged for self-referenced multiplex coherent anti-Stokes Raman scattering (SR-M-CARS) imaging [10, 5]. In particular, for imaging and other broader areas of application, there has recently been considerable interest in extending the length of a filament, or generating a ”superfilament” [11]. By enhancing our understanding of the dynamics of focusing-defocusing cycles in YAG and their concatenation, we envisage to design experiments which may produce longer filaments in the laboratory.
@@
Soliton Concatenation For a peak input power of 1.4 MW (equal to the critical power $P_{\text{cr}}$), a single self-focusing process appears, and spectral broadening accompanies the trapped spatial beam (Fig. 4a). After a few hundred microns of propagation the solitary wave spreads out, owing to the defocusing action of plasma generation and the emission of conical waves. By increasing the peak input power, a second self-focusing process appears, as shown in Fig. 4a. With a further peak input power increase, we observe that the second beam gradually approaches and concatenates with the first, thus substantially increasing the overall propagation length of the initially trapped wave (Fig.4a). For higher peak powers, the length of the solitary wave exceeds 5 mm, which is 5 times the filament length at 1.4 MW. The concatenation of Kerr-solitons and SCG necessary for SR-M-CARS is also represented by the numerical simulation results shown in Fig. 4b and Fig. 6a-c, respectively. Notably, a similar experiment carried out with a YAG crystal instead of Nd:YAG ceramics led to different dynamics (see Fig.5 and Visualizations 1 and 2 of supplementary material), which keep the individual trapped beams spatially separated . This suggests that a specially-conceived ceramic YAG with different dielectric properties could generate a more uniform and longer train of Townes solitons with enhanced robustness to the peak pump power. Plasma rate equation parameters such as the collision time $\tau_{c}$ (which is proportional to the electrical conductivity) influence the efficiency of plasma generation, particularly the magnitude of the plasma defocusing term in Eqn. 1, the maximum intensity (and minimum beam width), and the distance between focusing-defocusing cycles [1]. Therefore, the material parameter $\tau_{c}$ has an important effect on the maximum length of filaments.
@@
Conclusion In conclusion, we experimentally demonstrated that spatial solitons, generated at different positions in a ceramic polycrystal under the combined action of Kerr self-focusing and plasma defocusing, may merge and produce a continuous and stable laser filament, which propagates over several millimeters. We have presented a study of the defocusing action, filament onset location, SCG and conical emission in crystal and ceramic YAG. Moreover, we exploited the broadband output spectrum of the trapped beam with a clean spatial shape for SR-M-CARS imaging experiments. Optimizing the dielectric characteristics of ceramics may additionally make it possible to control the plasma defocusing process, and mitigate spatial instabilities of two-dimensional Kerr solitons. Further details appear in Section 2 of Ref. [19].

### v5

@@ Abstract
A coherent concatenation of multiple solitary waves may lead to a stable infrared and visible broadband filament in ceramic YAG polycrystal. This self-trapped soliton train is leveraged to implement self-referenced multiplex coherent anti-Stokes Raman scattering imaging. Simulations and experiments illustrating the filamentation process and the concatenation of focusing-defocusing cycles in ceramic and crystal YAG are presented. In addition, our simulations and experiments further examine the dependence of the filamentation onset location and supercontinuum generation upon peak input power. Understanding this dependence is key for implementation of viable CARS imaging techniques, due to the comparatively exceptional ability of YAG to generate supercontinuum which can enable higher-sensitivity imaging without delay lines.
Introduction Spatial solitons are formed by the deformation-free propagation of a wave packet, confined by the nonlinear effect of its own intensity [1]. Nonlinear saturation allows for wave stabilization over several Fresnel lengths, leading to a modulation-free, axially-symmetric beam. So-called Townes solitons can exactly balance diffraction via nonlinear effects at a given peak power. In two or more spatial dimensions, additional defocusing effects, such as plasma defocusing and nonlinear absorption, are known to arrest “catastrophic collapse” [1, 2, 3]. This process gives, at best, quasi-stability and can otherwise lead to cycles of focusing-defocusing events in which the Townes soliton is never entirely realized; instead, focusing-defocusing cycles of a solitary wave represent non-uniform oscillations around this Townes soliton state [3, 4, 5, 6, 7, 8]. It is also known that the extreme confinement of the field produces conical waves, causing some of the beam energy to propagate in a cone shape towards the center and feed the core filament due to a coherent interference effect [1, 5, 8].
In this work, we show experimentally and numerically that the self-focusing process obtained in a Nd:YAG ceramic rod can give rise to coherent, concatenated solitary waves, producing a stable filament in both visible and infrared spectra. This stabilization is obtained thanks to the particular dielectric properties of ceramics, which facilitate emission of conical waves, plasma defocusing, and nonlinear/plasma absorption [5, 9]. Such broadband supercontinuum (SC) is surprisingly stable, and can be efficiently leveraged for self-referenced multiplex coherent anti-Stokes Raman scattering (SR-M-CARS) imaging [10, 5]. In particular, for imaging and other broader areas of application, there has recently been considerable interest in extending the length of a filament, or generating a ”superfilament” [11]. By enhancing our understanding of the dynamics of focusing-defocusing cycles in YAG and their concatenation, we envisage to design experiments which may produce longer filaments in the laboratory.
@@
Soliton Concatenation For a peak input power of 1.4 MW (equal to the critical power $P_{\text{cr}}$), a single self-focusing process appears, and spectral broadening accompanies the trapped spatial beam (Fig. 4a). After a few hundred microns of propagation the solitary wave spreads out, owing to the defocusing action of plasma generation and the emission of conical waves. By increasing the peak input power, a second self-focusing process appears, as shown in Fig. 4a. With a further peak input power increase, we observe that the second beam gradually approaches and concatenates with the first, thus substantially increasing the overall propagation length of the initially trapped wave (Fig.4a). For higher peak powers, the length of the solitary wave exceeds 5 mm, which is 5 times the filament length at 1.4 MW. The concatenation of Kerr-solitons and SCG necessary for SR-M-CARS is also represented by the numerical simulation results shown in Fig. 4b and Fig. 6a-c, respectively. Notably, a similar experiment carried out with a YAG crystal instead of Nd:YAG ceramics led to different dynamics (see Fig.5 and Visualizations 1 and 2 of supplementary material), which keep the individual trapped beams spatially separated . This suggests that a specially-conceived ceramic YAG with different dielectric properties could generate a more uniform and longer train of solitary waves with enhanced robustness to the peak pump power. Plasma rate equation parameters such as the collision time $\tau_{c}$ (which is proportional to the electrical conductivity) influence the efficiency of plasma generation, particularly the magnitude of the plasma defocusing term in Eqn. 1, the maximum intensity (and minimum beam width), and the distance between focusing-defocusing cycles [1]. Therefore, the material parameter $\tau_{c}$ has an important effect on the maximum length of filaments.
@@
Conclusion In conclusion, we experimentally demonstrated that spatial solitons, generated at different positions in a ceramic polycrystal under the combined action of Kerr self-focusing and plasma defocusing, may merge and produce a continuous and stable laser filament, which propagates over several millimeters. We have presented a study of the defocusing action, filament onset location, SCG and conical emission in crystal and ceramic YAG. Moreover, we exploited the broadband output spectrum of the trapped beam with a clean spatial shape for SR-M-CARS imaging experiments. Optimizing the dielectric characteristics of ceramics may additionally make it possible to control the plasma defocusing process, and mitigate spatial instabilities of two-dimensional Kerr solitons.